OpenStudy (lemanual):

Help with 2 math questions please?

OpenStudy (lemanual):

http://prntscr.com/dfoyjf

OpenStudy (lemanual):

http://prntscr.com/dfoyqy

OpenStudy (mathstudent55):

In the first problem, each choice has a multiplication of two quantities. First, multiply the numbers together. For example, in A: \(\large 6a^5(6a^5) = 6 \times 6 \times a^5 \times a^5\)

OpenStudy (mathstudent55):

First, multiply \(6 \times 6\) What is that equal to?

OpenStudy (lemanual):

36

OpenStudy (mathstudent55):

Good. Now you need to multiply \(\large a^5 \times a^5\)

OpenStudy (lemanual):

a^10x?

OpenStudy (lemanual):

or 25

OpenStudy (mathstudent55):

To multiply powers with the same base, write the base and ADD the exponents. For example, \(\large b^5 \times b^7 = b^{5 + 7} = b^{12} \)

OpenStudy (mathstudent55):

Now try it with \(\large a^5 \times a^5\)

OpenStudy (lemanual):

a10

OpenStudy (mathstudent55):

Correct. That means that \(\large 6a^5 (6a^5) = 36a^{10} \) Is that the solution in choice A?

OpenStudy (lemanual):

ohhh so first and second?

OpenStudy (lemanual):

This? http://prntscr.com/dfp3v5

OpenStudy (mathmale):

Please use the symbol ^ to indicate exponentiation. correct: c^5 incorrect: c5

OpenStudy (lemanual):

O_O Yes sir

OpenStudy (mathstudent55):

Part of learning math (or anything else) is to pay attention. I just showed you what the correct answer to the problem in part A is. Now compare our correct answer with the answer in part A of your problem. Are the answers the same?

OpenStudy (mathstudent55):

We worked together and correctly on part A and got this. \(\large 6a^5 (6a^5) = 36a^{10}\) The problem has this: \(\large 6a^5 (6a^5) = 36a^{25}\) We did our solution correctly following the rules of multiplying powers. Since the problem has a different solution, and our solution is correct, the conclusion must be that part A of the problem is incorrect. We can even tell what their mistake is. They multiplied the exponents of variable a, when they should have added them.

OpenStudy (lemanual):

No so it's only the second one? cause you said add the exponents right? cause 6a^5+(6a^5)= 6x6=36 so thats correct but the exponent 5 you have to add and it says 25 which it isn't right?

OpenStudy (mathstudent55):

Remember, the questions asks which of the 4 parts were done CORRECTLY. Part A is not one of them.

OpenStudy (lemanual):

So do I multiply the exponents also?

OpenStudy (mathstudent55):

Have you read anything I wrote above?

OpenStudy (lemanual):

Yes I said you told me to add them but are you saying the first option is correct? cause if so how

OpenStudy (lemanual):

to add the exponents

OpenStudy (mathstudent55):

I just wrote this above: Remember, the questions asks which of the 4 parts were done CORRECTLY. Part A is not one of them.

OpenStudy (lemanual):

I added them and it's 5+5 which = 10 so that means it isn't first option

OpenStudy (mathstudent55):

Adding the exponents was correct for part A. We were correct. The problem is incorrect.

OpenStudy (lemanual):

Because it says 36a^25

OpenStudy (mathstudent55):

You are correct now. Option A is incorrect.

OpenStudy (lemanual):

Yeah so second option cause all of what I added was correct

OpenStudy (mathstudent55):

Option A of the problem, INCORRECTLY multiplied the powers when it should have added them.

OpenStudy (lemanual):

yeah I know so It's only second option correct?

OpenStudy (mathstudent55):

\(\large 5x^4(4x^2) = 5 \times 4 \times x^4 \times x^2 = 20x^6\) You are correct. Part B is correct.

OpenStudy (mathstudent55):

Now let's look at C.

OpenStudy (mathstudent55):

Here the numbers are positive and negative. What do you get when you multiply a positive number by a negative number?

OpenStudy (lemanual):

Positive number

OpenStudy (mathstudent55):

\(\large 6b^4(-3b^4) = 6 \times (-3) \times b^4 \times b^4\) What is 6 * (-3) = ?

OpenStudy (lemanual):

-18

OpenStudy (mathstudent55):

No. Rule for multiplying signed numbers: positive * positive = positive negative * negative = positive positive * negative = negative negative * positive = negative

OpenStudy (lemanual):

ohh

OpenStudy (mathstudent55):

Quick way of stating the rule: If you multiply two numbers with the same sign, the answer is positive. If you multiply two numbers with different signs, the answer is negative.

OpenStudy (mathstudent55):

Ok?

OpenStudy (lemanual):

Yes

OpenStudy (mathstudent55):

Good. Back to C. \(\large 6b^4(-3b^4) = 6 \times (-3) \times b^4 \times b^4\) Keep the rule above in mind. What is 6 * (-3) = ?

OpenStudy (mathstudent55):

6 is positive. -3 is negative. The numbers have different signs, so answer is? (positive or negative)

OpenStudy (lemanual):

Negative?

OpenStudy (mathstudent55):

Correct. What is 6 * 3?

OpenStudy (lemanual):

18

OpenStudy (mathstudent55):

Since 6 * 3 = 18, then 6 * (-3) = -18 Ok?

OpenStudy (lemanual):

I said -18 .-.

OpenStudy (mathstudent55):

Oh. I didn't see that. You did write positive above, and I was responding to that.

OpenStudy (mathstudent55):

Ok, so now we have -18. What is \(b^4 \times b^4\) ?

OpenStudy (lemanual):

8

OpenStudy (mathstudent55):

To multiply powers with the same base, we write the same base and ADD exponents.

OpenStudy (mathstudent55):

The exponent is 8. \(\large 6b^4(-3b^4) = 6 \times (-3) \times b^4 \times b^4 = -18b^8\)

OpenStudy (mathstudent55):

Is that the solution given in part C?

OpenStudy (lemanual):

Yes

OpenStudy (mathstudent55):

Good. Choose C also. Now you can do part D. What do you get for the product?

OpenStudy (lemanual):

I checked D and It's right because they added the exponent and got 9... 7+2=9 then 3*4 =12 and the answer is 12^9

OpenStudy (lemanual):

well 12z^9

OpenStudy (mathstudent55):

Excellent. You got the hang of this now. Good work!

OpenStudy (lemanual):

Yeah :D because of you :) ty

OpenStudy (mathstudent55):

Answer to the first problem, then, is B, C, D.

OpenStudy (mathstudent55):

Are you ready for the second problem now?

OpenStudy (lemanual):

Yes

OpenStudy (mathstudent55):

Ok. The second problem uses the first one as a building block. In the second problem, you need to use the distributive property in each part first. Once you have used the distributive property, then you multiply terms like we did in the first part.

OpenStudy (mathstudent55):

Are you familiar with the distributive property?

OpenStudy (lemanual):

I checked the first Option A and it's wrong :D

OpenStudy (lemanual):

should be 6 instead of 8 right?

OpenStudy (mathstudent55):

Here it is: The distributive property of multiplication over addition \(a(b + c) = ab + ac\)

OpenStudy (lemanual):

ohh yeah I know that

OpenStudy (lemanual):

I got the hang of it now :D can I do the rest and you just tell me if it's correct :))

OpenStudy (mathstudent55):

Now let's look at part A. \(\large -6y^4(4y^2 + 2)\) The first step is to apply the distributive property: \(\large -6y^4(4y^2 + 2) = -6y^4(4y^2) +~(- 6)y^4(2) =\) Now we do each multiplication as we did for the first problem. \(\large = -24y^6 = 12y^4\) You are correct. It is an exponent of 6, not 8.

OpenStudy (mathstudent55):

You learned well. Good job.

OpenStudy (lemanual):

http://prntscr.com/dfpkib ty so much I understand it now :)

OpenStudy (lemanual):

I am having trouble with this one though if you don't mind http://prntscr.com/dfpl9d

OpenStudy (mathstudent55):

Correct. Once again, great job!

OpenStudy (lemanual):

Because of you :) and ty

OpenStudy (lemanual):

Are you there?

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (lemanual):

nvm it was A :D

OpenStudy (lemanual):

for this one http://prntscr.com/dfpl9d

OpenStudy (lemanual):

Bye :) learn something new every day